# Spin-dependent Force from an NV center Ensemble on a Microlever

### v1

@@ II.3 Estimation of the NV centers magnetic torque
where $\gamma_{\rm e}=-(2\pi)28.0~{}{\rm GHz.T}^{-1}$ is the gyromagnetic factor of the NV center, $D$ is the anisotropy coefficient, $\hat{\mathbf{e}}_{3}$ is the anisotropy direction of the NV center (see the inset of Fig. 1-(a)), $\mathbf{B}(t)=\mathbf{B}_{0}+\mathbf{B}_{1}(t)$ is the sum of an homogeneous magnetic field and a time dependent microwave field. The magnetic field $\mathbf{B}_{0}$ is chosen not to be aligned with the crystalline anisotropy axis $\hat{\mathbf{e}}_{3}$, such that a magnetic torque may occur, as previously explained. The microwave field is here employed to drive an electronic spin resonance. We denote ={|$\mathcal{B}=\{\ket{0^{\prime}},\ket{-1^{\prime}},\ket{+1^{\prime}}\}$ as the basis of the new eigenstates of the NV center’s Hamiltonian in the magnetic field. Note that the magnetic moment may be non zero in the ground state |$\ket{0^{\prime}}$ because of its mixing with the |±$\ket{\pm 1}$ magnetic states. This effect can lead to magnetic torques even in the absence of microwaves Pellet-Mary et al. (2021); Perdriat (2023).
Dissipation must also be taken into account to calculate the magnetic moment of the NV centers. The local environment, namely paramagnetic impurities as well as lattice phonons, are responsible for a decay rate $\Gamma_{1}\approx(2\pi)1~{}{\rm kHz}$ and a dephasing rate $\Gamma_{2}^{*}\approx(2\pi)5~{}{\rm MHz}$ of the NV centers. Furthermore, the green laser polarizes the NV centers in the |$\ket{0^{\prime}}$ with a rate $\gamma_{\rm las}\approx(2\pi)1~{}{\rm kHz}$ in our experiment. This mechanism is efficient if the transverse magnetic field is smaller than 0.05 mT.
By driving a microwave resonant with the magnetic transition |$\ket{0^{\prime}}\to\ket{\pm 1^{\prime}}$, we can change the magnetic moment, thereby modifying the magnetic torque. The total magnetic moment variation is thus given by $\Delta\bm{\mu}=(\bm{\mu}_{\pm 1^{\prime}}-\bm{\mu}_{0^{\prime}})/2$, assuming an initial population in the |$\ket{0^{\prime}}$ state of $100\%$ and a final population of $50-50\%$ in |$\ket{0^{\prime}}$ and |±$\ket{\pm 1^{\prime}}$ states. Fig. 2-(b) shows a numerical estimation of the exerted magnetic torque as a function of the magnetic field strength. The magnetic field orientation has been chosen such that it makes an angle of $60^{\circ}$ with respect to the NV center axis. The magnetic torque is found to be on the order of $10^{-15}~{}{\rm N.m}$ at a magnetic field on the order of $10~{}{\rm mT}$. This value is three orders of magnitude above our torque sensitivity. Notably also, the theoretical torque exerted while driving the |$\ket{0^{\prime}}\to\ket{-1^{\prime}}$ transition should be stronger than that of the |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ transition.
@@ III.1 Detection of the NV center magnetic torque
Fig. 3-(a) shows an ODMR measurement (blue curve) of the diamond glued to the cantilever. We observe eight dips in the collected photoluminescence, corresponding to the |$\ket{0^{\prime}}\to\ket{-1^{\prime}}$ and |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ transitions of the four NV classes. Other small dips are visible and have been attributed to the presence of an unintended smaller diamond, likely added to the larger mono-crystalline $40~{}\mu{\rm m}$ diamond during the gluing process. We then scanned the microwave frequency while performing FM-MDMR in the same magnetic field configuration. The modulation amplitude was set to $8~{}{\rm MHz}$, which is on the order of the NV center linewidth. We measure the in-phase quadrature of the demodulated position signal from the interferometer (this point will be discussed in detail in Sec. III.3).
@@
The spin-torque measurements are presented in Fig. 3-(a) (red curve). Dispersive signals that all coincide with the ODMR dips are clearly observed, with cantilever displacements on the order of 0.2 nm for the central transitions. The dispersive curves are flipped in sign for the |$\ket{0^{\prime}}\to\ket{-1^{\prime}}$ and |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ resonances. This fact is more clearly observed in Fig. 3-(b) showing a zoom in on the two resonances of the same class at $2.88~{}{\rm GHz}$ and $3.00~{}{\rm GHz}$ respectively. Such opposite behavior indicates that the torque exerted by driving the two transitions are opposite, as expected. Furthermore, the torque exerted by driving the |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ transition is consistently smaller than the one on the |$\ket{0^{\prime}}\to\ket{-1^{\prime}}$ transition. This is in agreement with the theoretical expectations presented in Fig. 2-(b). One other possible explanation of the latter observation could be the microwave power dependency on the microwave frequency. The microwave signal power could indeed be lower in the high frequency range corresponding to the |$\ket{0^{\prime}}\to\ket{+1^{\prime}}$ transitions. If it was lower in that range however, the ODMR signals would also show a smaller contrast.
The torque measured on the cantilever is on the order of $\tau=k_{m}l_{\rm e}x_{\rm res}/Q=1.0\times 10^{-17}~{}{\rm N.m}$, where $x_{\rm res}=0.2~{}{\rm nm}$ is the amplitude of the cantilever displacement in FM-MDMR. The measured torque is two orders of magnitude lower than is predicted by the simulation shown in Fig. 2-(b). This discrepancy may be attributed to multiple factors. First, the green laser polarization efficiency is not $100\%$, as was assumed in the theoretical calculation. Second, the magnetic torque calculated in Fig. 2-(b) is the norm of the magnetic torque and not its projection along the $y$-axis. Third, the microwave power was not pushed to its maximum value in order to avoid adding spurious forces given to the cantilever motion. Fourth, we chose the in-phase quadrature of the displacement measurement. The in-phase quadrature is in fact not expected to carry any information about the resonator position if the torque were modulated in phase with the mechanical motion. In our case, however, the spin population does not react instantaneously to the frequency modulation, which give some signal to the X quadrature (see Sec. III.3). Finally, it has been previously observed that diamonds may have a density of NV centers lower than the specification provided by the company.
@@ III.2 Dependence of the magnetic torque on the magnetic field
Although many aspects point towards pure cross-product spin-torques as the origin of the cantilever displacement, it may sill arise from other effects. A possible cause of spin dependent cantilever displacement is a magnetic force. The magnetic field generated by the magnet is indeed not perfectly homogeneous, giving rise to magnetic field gradient on the order of $\nabla B\approx 1~{}{\rm T.m}^{-1}$, leading to a magnetic force of about $10^{-14}~{}{\rm N}$. This force is one order of magnitude lower than the calculated force arising from the magnetic torque, so it is expected to play a minor role in our experiment. One way to experimentally distinguish a force originating from pure magnetic torque from a force arising due to the magnetic field gradient is by varying the angle $\theta$ between the magnetic field and the NV center axis. A force due to magnetic torque would cancel out when $\theta=0$, whereas a force caused by the magnetic field gradient would remain unaffected. Fig. 4 (a)-(i) shows FM-MDMR measurements where the angle $\theta$ is varied while keeping the magnetic field at a constant value $B_{0}=18~{}{\rm mT}$. The angles are extracted by diagonalizing the Hamiltonian while varying the magnetic field amplitude and angle to match the energy of the pair of ODMR dips for one NV class. Notably, no mechanical displacement of the cantilever is observed for $\theta\approx 1^{\circ}$. As observed in Fig. 4 (a)-(ii), the torque indeed increases with $\theta$.
@@ III.3 NV center spin population dynamics
As previously mentioned, the green laser power ($1~{}{\rm mW}$) was kept low to avoid affecting the mechanical properties of the cantilever. As a result, the polarization rate is insufficient for the spin population to instantaneously follow the modulation of the microwave frequency. Consequently, the force applied to the mechanics is delayed relative to the frequency modulation excitation by approximately the inverse of the total spin polarization rate, $\gamma_{\rm tot}^{-1}=(\Gamma_{1}+\gamma_{\rm las}+\frac{\Omega^{2}}{2\Gamma_{2}^{*}})^{-1}$, which is on the order of milliseconds. This value is comparable to the period of the mechanical oscillator considered in our experiment. Therefore, the modulation of the microwave frequency at the mechanical resonance frequency excites both the X and Y quadratures, whereas only the Y quadrature would be expected to respond if the force were not delayed relative to the excitation.
@@
This delay in the spin population could actually be advantageous in mitigating spurious mechanical motion induced by the microwave signal. It has been observed that the microwave signal generates unwanted forces on the cantilever, which are primarily present in the Y quadrature, although the exact origin of these forces is not yet fully understood. These forces could be of magnetic, electric, or acoustic origin, potentially stemming from acoustic noise caused by heating of the microwave wire, or from magnetic or electric field gradients generated by the microwave, which exert forces on the slightly diamagnetic/dielectric cantilever. By focusing on the X quadrature, one can thus suppress noise originating from the microwave that is in phase with the microwave frequency modulation.
@@ IV Conclusion
In conclusion, we have observed a force induced by the spin of negatively charged NV centers on a clamped micro-mechanical oscillator. This was achieved by frequency modulating the microwave drive at resonance with the cantilever’s fundamental flexural mode in the presence of a homogeneous magnetic field. The angular dependence of the force on the direction of the homogeneous magnetic field indicates that the spin torque, via the cross-product, is responsible for coupling to the flexural mode. Homogeneous magnetic fields can be advantageous over gradients, especially when working with spin ensembles larger than a few $\mu$m³, where the inhomogeneous broadening of the spin ensemble becomes excessive due to strong field gradients.
@@ Acknowledgements.
Appendix A Diamond positioning
To attach the diamond to the cantilever, the process starts by applying a drop of glue to the tip of a metallic wire with a diameter of 15 µm. Using a 3D micrometric translation stage, the wire is carefully moved so that the glue touches the end of the cantilever. Next, a single diamond is held by Van der Waals forces at the end of another 15 µm wire. With the same micrometric stage, the diamond is gently placed onto the glued end of the cantilever. All of these steps are performed under an optical microscope. The glue is then allowed to dry for a few minutes, securing the diamond in place.
Appendix B Mode profile of the first bending mode of the cantilever
@@
Appendix C Estimation of the magnetic moment of the NV center’s eigenstates
Figure 7: (a) Single spin magnetic moments in the three NV eigenstates as a function of magnetic field with an amplitude ranging from $0.01$ to $0.3$ T. (b) Magnetic torque in the three eigenstates as a function of magnetic field strength, ranging from 0 to 0.3 T.
Fig. 7 (a) shows a calculation of single spin magnetic moments of an NV center. It is calculated in the three NV eigenstates, for different values of the magnetic field strength $B_{0}$ in a 2D plane formed by $\mathbf{B}_{0}$ and the NV axis $\mathbf{e}_{3}$. We assume here that the magnetic field orientation is kept at $60^{\circ}$ with respect to the anisotropy axis. In the small magnetic field regime ($\gamma_{\rm e}B_{0}\ll D$), the eigenstates magnetic moment orientation is oriented along the anisotropy axis, and does not depend on the magnetic field orientation. In the strong magnetic field case ($\gamma_{\rm e}B_{0}\gg D$), the |$\ket{+1^{\prime}}$ state tends to align to the magnetic field direction, the |$\ket{0^{\prime}}$ state tends to align oppositely to the magnetic field direction and the |$\ket{-1^{\prime}}$ state tends to a null magnetization. In the range $\gamma_{\rm e}B_{0}\approx D$, the magnetic moment evolves continuously between theses two extreme cases. The evolution of the magnetic moment orientation and intensity as a function of the magnetic field directly impacts the resulting magnetic torque in the three eigenstates.
Fig. 7 (b) shows the result of numerical calculations of the magnetic torque amplitude for a single spin, in each NV eigenstates. We assume here $100\%$ population in each single eigenstate, while keeping the same magnetic field orientation. At small magnetic field values, the magnetic torque in the |±$\ket{\pm 1^{\prime}}$ follows a linear trend, while the torque in the |$\ket{0^{\prime}}$ state is quadratic. For strong magnetic fields, the magnetic torque amplitude in the |$\ket{0^{\prime}}$ and |$\ket{+1^{\prime}}$ states converges towards a value that is half the opposite of the torque in the |$\ket{-1^{\prime}}$ state. It is on the order of $10^{-14}$ N.m. Surprisingly, the magnetic torque converges towards a value that is non-zero even if the magnetization in each eigenstates tends to align with the magnetic field or tends to zero. This is due to the fact that the magnetization has a contribution perpendicular to the magnetic field that scales as $D/B_{0}$. At large magnetic fields, the magnetic torque amplitude is thus on the order of $\hbar ND$. In practice, the polarization efficiency of the green laser into the |$\ket{0^{\prime}}$ state is no more efficient at magnetic field values in the $0.1$ T range.
Appendix D Calculation of the NV center torques
@@
where $\hat{\mathbf{L}}$ is the angular momentum operator of the diamond, $I$ the moment of inertia (the diamond is assumed to be spherical), $D=(2\pi)2.87~{}{\rm GHz}$ is the crystalline anisotropy of the NV center, $\hat{\mathbf{S}}^{(i)}$ is the spin-1 operator of the i-th NV center in the diamond, $N$ is the number of NV centers, $\hat{\mathbf{e}}_{3}$ is the crystalline anisotropy orientation normalized vector, and $\mathbf{B}(t)=\mathbf{B}_{0}+\mathbf{B}_{1}(t)$ is the sum of the homogeneous and microwave magnetic field.
@@
$\frac{\mathrm{d\mathbf{L}}}{\mathrm{d}t}$ $=\frac{1}{i\hbar}\left\langle\commutator{\hat{\mathbf{L}}}{\hat{\mathcal{H}}_{\rm s.m.}}\right\rangle,$ ] ⟩ , (8)
$\forall i\in[1,N],\quad\frac{\mathrm{d}\mathbf{S}^{(i)}}{\mathrm{d}t}$ $=\frac{1}{i\hbar}\left\langle\commutator{\hat{\mathbf{S}}^{(i)}}{\hat{\mathcal{H}}_{\rm s.m.}}\right\rangle.$ ] ⟩ . (9)
The angular momentum operator $\hat{\mathbf{L}}$ obviously commutes with $\hat{\mathbf{L}}^{2}$ and the spin operators $\hat{\mathbf{S}}^{(i)}$, which exist in a different Hilbert space. The crystalline anisotropy direction $\hat{\mathbf{e}}_{3}$ depends on the angular position of the diamond, given by the three Euler angles $(\hat{\alpha},\hat{\beta},\hat{\gamma})$, which do not commute with the angular momentum $\hat{\mathbf{L}}$. This results in a torque exerted by the NV centers on the mechanics. Furthermore, the spin observable $\hat{\mathbf{S}}^{(i)}$ commutes with $\hat{\mathbf{L}}^{2}$, but it does not commute with the projection of the spin operator onto the vector observable $\hat{\mathbf{e}}_{3}$ or onto the magnetic field $\mathbf{B}(t)$. These commutators can be calculated (see Appendix B in Perdriat (2023)) and lead to the following equations:
@@
The first torque, $\bm{\mu}\times\mathbf{B}_{0}$, represents the magnetic torque applied to the diamond due to the homogeneous magnetic field. The second torque $\bm{\mu}\times\mathbf{B}_{1}(t)$ is the magnetic torque applied to the diamond from the microwave field. The microwave field oscillates at gigahertz frequencies, so this term could be expected to average out to zero. However, while the magnetic resonance is driven, the spin magnetic moment rotates perpendicular to the microwave field, resulting in a non-zero average magnetic torque. This term is the classical equivalent of the torque resulting from the absorption of circularly polarized microwave photons. Finally, the last torque $-1/\gamma_{\rm e}\mathrm{d}\bm{\mu}/\mathrm{d}t$ corresponds to the Einstein-de Haas torque.
Appendix E Calculation of the mechanical response to the modulation of the NV center population
In this section, we determine the mechanical response of the cantilever to the frequency modulation of the microwave. For simplicity, we assume that the resonance |$\ket{0^{\prime}}\to\ket{\pm 1^{\prime}}$ is driven on the blue side of the transition. This means that the microwave detuning is chosen such that $\Delta(t)=\Gamma_{2}^{*}+\delta\omega_{\mu}(t)$, where $\delta\omega_{\mu}(t)\ll\Gamma_{2}^{*}$ represents the frequency modulation, which is considered small compared to the decoherence rate. This assumption is not really verified in our experiment since $\delta\omega_{\mu}(t)$ is on the order of $\Gamma_{2}^{*}$, but is done to simplify the calculation.

### v2

@@ II.3 Estimation of the NV centers magnetic torque
where $\gamma_{\rm e}=-(2\pi)28.0\leavevmode\nobreak\ {\rm GHz.T}^{-1}$ is the gyromagnetic factor of the NV center, $D$ is the anisotropy coefficient, $\hat{\mathbf{e}}_{3}$ is the anisotropy direction of the NV center (see the inset of Fig. 1-(a)), $\mathbf{B}(t)=\mathbf{B}_{0}+\mathbf{B}_{1}(t)$ is the sum of an homogeneous magnetic field and a time dependent microwave field. The magnetic field $\mathbf{B}_{0}$ is chosen not to be aligned with the crystalline anisotropy axis $\hat{\mathbf{e}}_{3}$, such that a magnetic torque may occur, as previously explained. The microwave field is here employed to drive an electronic spin resonance. We denote =$\mathcal{B}=\{|{0^{\prime}}\rangle,|{-1^{\prime}}\rangle,|{+1^{\prime}}\rangle\}$ as the basis of the new eigenstates of the NV center’s Hamiltonian in the magnetic field. Note that the magnetic moment may be non zero in the ground state $|{0^{\prime}}\rangle$ because of its mixing with the |±$|{\pm 1}\rangle$ magnetic states. This effect can lead to magnetic torques even in the absence of microwaves Pellet-Mary et al. (2021); Perdriat (2023).
Dissipation must also be taken into account to calculate the magnetic moment of the NV centers. The local environment, namely paramagnetic impurities as well as lattice phonons, are responsible for a decay rate $\Gamma_{1}\approx(2\pi)1\leavevmode\nobreak\ {\rm kHz}$ and a dephasing rate $\Gamma_{2}^{*}\approx(2\pi)5\leavevmode\nobreak\ {\rm MHz}$ of the NV centers. Furthermore, the green laser polarizes the NV centers in the $|{0^{\prime}}\rangle$ with a rate $\gamma_{\rm las}\approx(2\pi)1\leavevmode\nobreak\ {\rm kHz}$ in our experiment. This mechanism is efficient if the transverse magnetic field is smaller than 0.05 mT.
By driving a microwave resonant with the magnetic transition $|{0^{\prime}}\rangle\to|{\pm 1^{\prime}}\rangle$, we can change the magnetic moment, thereby modifying the magnetic torque. The total magnetic moment variation is thus given by $\Delta\bm{\mu}=(\bm{\mu}_{\pm 1^{\prime}}-\bm{\mu}_{0^{\prime}})/2$, assuming an initial population in the $|{0^{\prime}}\rangle$ state of $100\%$ and a final population of $50-50\%$ in $|{0^{\prime}}\rangle$ and |±$|{\pm 1^{\prime}}\rangle$ states. Fig. 2-(b) shows a numerical estimation of the exerted magnetic torque as a function of the magnetic field strength. The magnetic field orientation has been chosen such that it makes an angle of $60^{\circ}$ with respect to the NV center axis. The magnetic torque is found to be on the order of $10^{-15}\leavevmode\nobreak\ {\rm N\cdot m}$ at a magnetic field on the order of $10\leavevmode\nobreak\ {\rm mT}$. This value is three orders of magnitude above our torque sensitivity. Notably also, the theoretical torque exerted while driving the $|{0^{\prime}}\rangle\to|{-1^{\prime}}\rangle$ transition should be stronger than that of the $|{0^{\prime}}\rangle$ to $|{+1^{\prime}}\rangle$ transition.
@@ III.1 Detection of the NV center magnetic torque
Fig. 3-(a) shows an ODMR measurement (blue curve) of the diamond glued to the cantilever. We observe eight dips in the collected photoluminescence, corresponding to the $|{0^{\prime}}\rangle\to|{-1^{\prime}}\rangle$ and $|{0^{\prime}}\rangle\to|{+1^{\prime}}\rangle$ transitions of the four NV classes. Other small dips are visible and have been attributed to the presence of an unintended smaller diamond, likely added to the larger mono-crystalline $40\leavevmode\nobreak\ \mu{\rm m}$ diamond during the gluing process. We then scanned the microwave frequency while performing FM-MDMR in the same magnetic field configuration. The modulation amplitude was set to $8\leavevmode\nobreak\ {\rm MHz}$, which is on the order of the NV center linewidth. We measure the in-phase quadrature of the demodulated position signal from the interferometer (this point will be discussed in detail in III.3).
@@
The spin-torque measurements are presented in Fig. 3-(a) (red curve). Dispersive signals that all coincide with the ODMR dips are clearly observed, with cantilever displacements on the order of 0.2 nm for the central transitions. The dispersive curves are flipped in sign for the $|{0^{\prime}}\rangle\to|{-1^{\prime}}\rangle$ and $|{0^{\prime}}\rangle\to|{+1^{\prime}}\rangle$ resonances. This fact is more clearly observed in Fig. 3-(b) showing a zoom in on the two resonances of the same class at $2.88\leavevmode\nobreak\ {\rm GHz}$ and $3.00\leavevmode\nobreak\ {\rm GHz}$ respectively. Such opposite behavior indicates that the torque exerted by driving the two transitions are opposite, as expected. Furthermore, the torque exerted by driving the $|{0^{\prime}}\rangle\to|{+1^{\prime}}\rangle$ transition is consistently smaller than the one on the $|{0^{\prime}}\rangle\to|{-1^{\prime}}\rangle$ transition. This is in agreement with the theoretical expectations presented in Fig. 2-(b). One other possible explanation of the latter observation could be the microwave power dependency on the microwave frequency. The microwave signal power could indeed be lower in the high frequency range corresponding to the $|{0^{\prime}}\rangle\to|{+1^{\prime}}\rangle$ transitions. If it was lower in that range however, the ODMR signals would also show a smaller contrast.
The torque measured on the cantilever is of the order of $\tau=k_{m}l_{\rm e}x_{\rm res}/Q=1.0\times 10^{-17}\leavevmode\nobreak\ {\rm N\cdot m}$, where $x_{\rm res}=0.2\leavevmode\nobreak\ {\rm nm}$ is the amplitude of the cantilever displacement in FM-MDMR. The measured torque is two orders of magnitude lower than is predicted by the simulation shown in Fig. 2-(b). This discrepancy may be attributed to multiple factors. First, the green laser polarization efficiency is not $100\%$, as was assumed in the theoretical calculation. Second, the magnetic torque calculated in Fig. 2-(b) is the norm of the magnetic torque and not its projection along the $y$-axis. Third, the microwave power was not pushed to its maximum value in order to avoid adding spurious forces given to the cantilever motion. Fourth, we chose the in-phase quadrature of the displacement measurement. The in-phase quadrature is in fact not expected to carry any information about the resonator position if the torque were modulated in phase with the mechanical motion. In our case, however, the spin population does not react instantaneously to the frequency modulation, which give some signal to the X quadrature (see III.3). Finally, it has been previously observed that diamonds may have a density of NV centers lower than the specification provided by the company.
@@ III.2 Dependence of the magnetic torque on the magnetic field
Although many aspects point towards pure cross-product spin-torques as the origin of the cantilever displacement, it may sill arise from other effects. A possible cause of spin dependent cantilever displacement is a magnetic force. The magnetic field generated by the magnet is indeed not perfectly homogeneous, giving rise to magnetic field gradient of the order of $\nabla B\approx 1\leavevmode\nobreak\ {\rm T\cdot m}^{-1}$, leading to a magnetic force of about $10^{-14}\leavevmode\nobreak\ {\rm N}$. This force is one order of magnitude lower than the calculated force arising from the magnetic torque, so it is expected to play a minor role in our experiment. One way to experimentally distinguish a force originating from pure magnetic torque from a force arising due to the magnetic field gradient is by varying the angle $\theta$ between the magnetic field and the NV center axis. A force due to magnetic torque would cancel out when $\theta=0$, whereas a force caused by the magnetic field gradient would remain unaffected. Fig. 4 (a)-(i) shows FM-MDMR measurements where the angle $\theta$ is varied while keeping the magnetic field at a constant value $B_{0}=18\leavevmode\nobreak\ {\rm mT}$. The angles are extracted by diagonalizing the Hamiltonian while varying the magnetic field amplitude and angle to match the energy of the pair of ODMR dips for one NV class. Notably, no mechanical displacement of the cantilever is observed for $\theta\approx 1^{\circ}$. As observed in Fig. 4 (a)-(ii), the torque indeed increases with $\theta$.
@@ III.3 NV center spin population dynamics
As previously mentioned, the green laser power ($1\leavevmode\nobreak\ {\rm mW}$) was kept low to avoid affecting the mechanical properties of the cantilever. As a result, the polarization rate is insufficient for the spin population to instantaneously follow the modulation of the microwave frequency. Consequently, the force applied to the mechanics is delayed relative to the frequency modulation excitation by approximately the inverse of the total spin polarization rate, $\gamma_{\rm tot}^{-1}=(\Gamma_{1}+\gamma_{\rm las}+\frac{\Omega^{2}}{2\Gamma_{2}^{*}})^{-1}$, which is of the order of milliseconds. This value is comparable to the period of the mechanical oscillator considered in our experiment. Therefore, the modulation of the microwave frequency at the mechanical resonance frequency excites both the X and Y quadratures, whereas only the Y quadrature would be expected to respond if the force were not delayed relative to the excitation.
@@
This delay in the spin population could actually be advantageous in mitigating spurious mechanical motion induced by the microwave signal. It has been observed that the microwave signal generates unwanted forces on the cantilever, which are primarily present in the Y quadrature, although the exact origin of these forces is not yet fully understood. These forces could be of magnetic, electric, or acoustic origin, potentially stemming from acoustic noise caused by heating of the microwave wire, or from magnetic or electric field gradients generated by the microwave, which exert forces on the slightly diamagnetic/dielectric cantilever. Focusing on the X quadrature, suppresses noise originating from the microwave that is in phase with the microwave frequency modulation.
@@ IV Conclusion
In conclusion, we have observed a force induced by the spin of negatively charged NV centers on a clamped micro-mechanical oscillator. This was achieved by frequency modulating the microwave drive at resonance with the cantilever’s fundamental flexural mode in the presence of a homogeneous magnetic field. The angular dependence of the force on the direction of the homogeneous magnetic field indicates that the spin torque, via the cross-product, is responsible for coupling to the flexural mode. Homogeneous magnetic fields can be advantageous over gradients, especially when working with spin ensembles larger than a few $\mu$m3, where the inhomogeneous broadening of the spin ensemble becomes excessive due to strong field gradients.
@@ Acknowledgements.
Annexe A Diamond positioning
To attach the diamond to the cantilever, the process starts by applying a drop of glue to the tip of a metallic wire with a diameter of 15 $\mu$m. Using a 3D micro-metric translation stage, the wire is carefully moved so that the glue touches the end of the cantilever. Next, a single diamond is held by Van der Waals forces at the end of another 15 $\mu$m wire. With the same micrometric stage, the diamond is gently placed onto the glued end of the cantilever. All of these steps are performed under an optical microscope. The glue is then allowed to dry for a few minutes, securing the diamond in place.
Annexe B Mode profile of the first bending mode of the cantilever
@@
Annexe C Estimation of the magnetic moment of the NV center’s eigenstates
Figure 7: (a) Single spin magnetic moments in the three NV eigenstates as a function of magnetic field with an amplitude ranging from 0.010010.010.01 to 0.3030.30.3 T, at an angle close to $\pi/4$ with respect to the NV axis as indicated in the figure. (b) Magnetic torque in the three eigenstates as a function of magnetic field strength, ranging from 0 to 0.3 T.
Fig. 7(a) presents a calculation of the single-spin magnetic moments of an NV center in its three eigenstates, for different values of the magnetic field strength $B_{0}$. The calculations are performed in a 2D plane defined by $\mathbf{B}_{0}$ and the NV axis $\mathbf{e}_{3}$, assuming that the magnetic field orientation remains fixed at $60^{\circ}$ relative to the anisotropy axis.
In the low-field regime ($\gamma_{\rm e}B_{0}\ll D$), the magnetic moment is aligned along the anisotropy axis and remains independent of the magnetic field orientation. In contrast, in the high-field regime ($\gamma_{\rm e}B_{0}\gg D$), the $|{+1^{\prime}}\rangle$ state aligns with the magnetic field direction, the $|{0^{\prime}}\rangle$ state aligns in the opposite direction, and the $|{-1^{\prime}}\rangle$ state exhibits nearly zero magnetization. In the intermediate regime ($\gamma_{\rm e}B_{0}\approx D$), the magnetic moment continuously evolves between these two limiting cases.
The variation in both the orientation and magnitude of the magnetic moment as a function of the magnetic field directly influences the resulting magnetic torque in each of the three eigenstates. Fig. 7 (b) shows the result of numerical calculations of the magnetic torque amplitude for a single spin, in each NV eigenstates. We assume here $100\%$ population in each single eigenstate, while keeping the same magnetic field orientation. At low magnetic field values, the magnetic torque in the |±$|{\pm 1^{\prime}}\rangle$ states follows a linear trend, whereas in the $|{0^{\prime}}\rangle$ state, it exhibits a quadratic dependence. For strong magnetic fields, the magnetic torque amplitude in the $|{0^{\prime}}\rangle$ and $|{+1^{\prime}}\rangle$ states converges to a value that is half the opposite of the torque in the $|{-1^{\prime}}\rangle$ state, of the order of $10^{-14}$ N·m.
Interestingly, the magnetic torque remains nonzero even when the magnetization in each eigenstate tends to align with the magnetic field or vanishes. This arises from the fact that the magnetization retains a perpendicular component to the magnetic field, which scales as $D/B_{0}$. Consequently, at large magnetic fields, the magnetic torque amplitude is of the order of $\hbar ND$. In practice, the polarization efficiency of the green laser into the $|{0^{\prime}}\rangle$ state becomes ineffective for magnetic field values around 0.1010.10.1 T.
Annexe D Calculation of the NV center torques
@@
where $\hat{\mathbf{L}}$ is the angular momentum operator of the diamond, $I$ the moment of inertia (the diamond is assumed to be spherical), $D=(2\pi)2.87\leavevmode\nobreak\ {\rm GHz}$ is the crystalline anisotropy of the NV center, $\hat{\mathbf{S}}^{(i)}$ is the spin-1 operator of the i-th NV center in the diamond, $N$ is the number of NV centers, $\hat{\mathbf{e}}_{3}$ is the crystalline anisotropy orientation normalized vector, and $\mathbf{B}(t)=\mathbf{B}_{0}+\mathbf{B}_{1}(t)$ is the sum of the homogeneous and microwave magnetic field.
@@
$\frac{\mathrm{d\mathbf{L}}}{\mathrm{d}t}$ $=\frac{1}{i\hbar}\left\langle[{\hat{\mathbf{L}}},{\hat{\mathcal{H}}_{\rm s.m.}}]\right\rangle,$ (8)
$\forall i\in[1,N],\quad\frac{\mathrm{d}\mathbf{S}^{(i)}}{\mathrm{d}t}$ $=\frac{1}{i\hbar}\left\langle[{\hat{\mathbf{S}}^{(i)}},{\hat{\mathcal{H}}_{\rm s.m.}}]\right\rangle.$ (9)
The angular momentum operator $\hat{\mathbf{L}}$ commutes with $\hat{\mathbf{L}}^{2}$ and the spin operators $\hat{\mathbf{S}}^{(i)}$. The crystalline anisotropy direction $\hat{\mathbf{e}}_{3}$ depends on the angular position of the diamond, given by the three Euler angles $(\hat{\alpha},\hat{\beta},\hat{\gamma})$, which do not commute with the angular momentum $\hat{\mathbf{L}}$. This results in a torque exerted by the NV centers on the mechanics. Furthermore, the spin observable $\hat{\mathbf{S}}^{(i)}$ commutes with $\hat{\mathbf{L}}^{2}$, but it does not commute with the projection of the spin operator onto the vector observable $\hat{\mathbf{e}}_{3}$ or onto the magnetic field $\mathbf{B}(t)$. These commutators can be calculated (see Appendix B in Perdriat (2023)) and lead to the following equations:
@@
The first torque, $\bm{\mu}\times\mathbf{B}_{0}$, represents the magnetic torque applied to the diamond due to the homogeneous magnetic field. The second torque $\bm{\mu}\times\mathbf{B}_{1}(t)$ is the magnetic torque applied to the diamond from the microwave field. The microwave field oscillates at GHz frequencies, so this term could be expected to average out to zero. However, while the magnetic resonance is driven, the spin magnetic moment rotates perpendicular to the microwave field, resulting in a non-zero average magnetic torque. This term is the classical equivalent of the torque resulting from the absorption of circularly polarized microwave photons. Finally, the last torque $-1/\gamma_{\rm e}\mathrm{d}\bm{\mu}/\mathrm{d}t$ corresponds to the Einstein-de Haas torque.
Annexe E Calculation of the mechanical response to the modulation of the NV center population
In this section, we determine the mechanical response of the cantilever to the frequency modulation of the microwave. For simplicity, we assume that the $|{0^{\prime}}\rangle\to|{\pm 1^{\prime}}\rangle$ resonance is driven on the blue side of the transition. This implies that the microwave detuning is chosen such that
$\Delta(t)=\Gamma_{2}^{*}+\delta\omega_{\mu}(t),$
where $\delta\omega_{\mu}(t)\ll\Gamma_{2}^{*}$ represents the frequency modulation, which is considered small compared to the decoherence rate. While this assumption does not hold in our experiment—since $\delta\omega_{\mu}(t)$ is of the order of $\Gamma_{2}^{*}$—it is made here to simplify the calculations.